# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

### v2

@@ Abstract.
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet measuring the ‘success’ of governance is complex and nuanced, making it challenging to determine what governance strategies are most successful. Furthermore, prior work has shown that communities have differing values, suggesting that ‘one-size-fits-all’ approaches to governance are unlikely to serve all communities well. In this work, we assess governance practices on reddit by classifying the sentiment of community members’ public discussion of their own moderators. We label 1.89 million posts and comments made on reddit over an 18 month period. We relate these perceptions to characteristics of community governance, and to different actions that community moderators take. We identify types of communities where moderators are perceived particularly positively and negatively, and highlight promising strategies for moderator teams. Amongst other findings, we show that strict rule enforcement is linked to more favorable perceptions of moderators of communities dedicated to certain topics, such as news communities, than others. We investigate what kinds of moderators are associated with improved community perceptions upon their addition to a mod team, and find that moderators who are active community members before and during their mod tenures are seen more favorably. We make all our models, datasets, and code public.
@@ 1. Introduction
We discuss the limitations of our methods, and identify key areas for future work (§7). We make our models, datasets, and code public [^1] Contact the first author for more information. to enable further research on this important topic.
@@ 2. Related Work
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera et al., 2009) and some gaming servers (Frey and Sumner, 2019). On reddit and most popular social media platforms, however access is difficult to obtain, and, and requires separate permission from the mods of every community to be studied (Li et al., 2022a, b). In contrast, our method works for every community whose content is public.
@@ 3. Methods
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al., 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, membership, and moderators.
In this work we make extensive use of publicly available reddit data from the Pushshift (Baumgartner et al., 2020). From these data, we detect and classify posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). We collect supplementary data about moderators (§3.1) and communities (§3.2).
@@ 3.1. Computing Moderator Timelines
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of the 10,000 largest subreddit’s moderators using the Wayback Machine, a web archiving service provided by the Internet Archive (Archive, 2024). We collect archived snapshots of every subreddit’s moderator info page going back to 2010, totalling 30,302 snapshots, which we combine with 10,000 present day snapshots collected directly from reddit. Each snapshot includes a list of moderators’ usernames, ordered by seniority. As the Wayback Machine snapshots higher traffic pages more frequently, we have more higher temporal resolution for larger and more popular subreddits. However, as each moderator info page contains the exact timestamp that each moderator was added, we have second-level precision for new moderator appointments. After a final data-cleaning step to remove known bots, the resulting dataset of moderator timelines consists of 89,237 moderator tenures across 9,502 subreddits.
@@ 3.2. Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the 10,000 largest subreddits, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld et al., 2022). We used the manually-labeled dataset of 123 subreddits from Weld et al. (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld et al., 2022). We average across all responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
@@ 3.3. Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying posts and comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify content written by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects content discussing moderators (differentiating them from those which make other use of ‘mod’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of this content into positive, neutral, negative, and exclude classes. We only include content written by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not on how mods discuss themselves in public. We apply our pipeline, which we make public, to all reddit posts and comments made from Jan. 2020 to June 2021.
@@
Detection Step Details (Step 2). To differentiate posts and comments which discuss moderators from those which use ‘mod(s),’ we fine-tuned a RoBERTa-based binary classifier (Liu et al., 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the entire reddit dataset and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa (Figure 1). To label posts and comments for the classification task, two annotators worked together to iteratively refine a codebook (Supplementary Materials A), then independently labeled 200 randomly sampled posts/comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch, 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts/comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts/comments for use as a training set, of which 484 were randomly sampled, and 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa model trained for the classification task.
@@
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators. For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with content that erroneously passed the detection step. We trained on this ‘exclude’ class in addition to those used for downstream analyses: positive, neutral, and negative.
To further enhance the performance of our model, we performed data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al., 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al., 2023) fine-tuned with QLoRA (Dettmers et al., 2023). LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step left us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make this dataset public.
@@
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4 & §6: The Amount of Mod Discourse for a given community is the fraction of all posts and comments discussing mods in that community, regardless of their sentiment. The Composition of Mod Discourse with positive/negative sentiment is the percent of all posts and comments discussing mods in a given community that have positive/negative sentiment. To ensure a meaningful amount of data, we exclude communities that did not have at least one positive, neutral, and negative post or comment discussing moderators. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. This filtering leaves us with 5,282 subreddits used in downstream analyses.
@@ 4. How are moderators of different communities perceived differently by their communities?
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities have more mod discourse (Figure 3a), with tiny communities with $<10$ posts+comments/day having 7.6×7.6$\times$7.6 × as many posts and comments discussing moderators as huge communities (those with more than 10k posts and comments per day), relative to the total amount of content. Tiny communities also have 6.0×6.0$\times$6.0 × more positive mod discourse, and 0.46×0.46$\times$0.46 × as much negative discourse, than huge communities.
@@ 5. Adjusting for Confounders using IPTW & DID
While we are ideally interested in causal, actionable insights, our study is a retrospective observational study, and like any other, subject to potential confounding. In §4 we identified two important confounders, size and topic. Next we describe two methods from the causal inference literature which we use to adjust for these and other confounding factors in subsequent analyses.
Inverse Probability of Treatment Weighting (IPTW). IPTW is a statistical method to adjust for imbalance in possible confounding factors when making comparisons between different observed groups (e.g., a treatment and a control group). We select IPTW over other methods, such as stratification, due to its excellent efficiency (Austin and Stuart, 2015). In §6, we compute the probability of treatment (propensity score, $P(Z|\mathbf{X})$) by applying a logistic regression model to community covariates $\mathbf{X}$ including the topic and size of the community. A complete list of covariates for each analysis is given in Supplementary Materials  C. We generalize IPTW to non-binary treatments to estimate a dose-response curve using the method from (Althoff et al., 2016) in which the treatment is discretized into bins and a different propensity score is computed for each treatment bin in a ‘one-versus-rest’ scheme. For final analyses, each observation is weighted by the inverse of the probability of the treatment it received, such that the weight for observation $i$ is $w_{i}=\frac{Z_{i}}{P(Z_{i}=1|\mathbf{X}_{i})}+\frac{1-Z_{i}}{P(Z_{i}=0|\mathbf{X}_{i})}$, where $Z_{i}$ is the treatment received by observation $i$ (0 for control, 1 for treated), and $\mathbf{X}_{i}$ is a vector of the covariates. In analyses where we use IPTW, we also include the non-IPTW adjusted estimate in gray. In these analyses, reweighting does not dramatically impact our findings, suggesting that the impact of these confounders is moderate at best.
An important validity check for IPTW is to assess the balance of covariates for each treatment group after weighting as applied (Austin and Stuart, 2015). Two groups are often considered ‘balanced’ or ‘indistinguishable‘ if all covariates are within a standardized mean difference (SMD) of 0.25 standard deviations (Althoff et al., 2016). For each treatment group, we compute the difference between each covariate’s weighted mean value and the reference distribution, consisting of the entire population. We compute the standardized mean difference (SMD) by normalizing the difference in means by the standard deviation of the values of the reference distribution. Our reweighting method achieves balance for every covariate substantially related to the treatment. SMDs after weighting are given for each covariate and each analysis in Supplementary Materials C.

### v3

@@ Abstract.
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet measuring the ‘success’ of governance is complex and nuanced, making it challenging to determine what governance strategies are most successful. Furthermore, prior work has shown that communities have differing values, suggesting that ‘one-size-fits-all’ approaches to governance are unlikely to serve all communities well. In this work, we assess governance practices on reddit by classifying the sentiment of community members’ public discussion of their own moderators. We label 1.89 million posts and comments made on reddit over an 18 month period. We relate these perceptions to characteristics of community governance, and to different actions that community moderators take. We identify types of communities where moderators are perceived particularly positively and negatively, and highlight promising strategies for moderator teams. Amongst other findings, we show that strict rule enforcement is linked to more favorable perceptions of moderators of communities dedicated to certain topics, such as news communities, than others. We investigate what kinds of moderators are associated with improved community perceptions upon their addition to a mod team, and find that moderators who are active community members before and during their mod tenures are seen more favorably. We make our models, anonymized datasets, and code public.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Social network analysisccs: Information systems Social networking sites
@@ 1. Introduction
We discuss the limitations of our methods, and identify key areas for future work (§7). We make our models, datasets, and code public [^1] Please contact the authors for additional information. to enable further research on this important topic.
@@ 2. Related Work
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera et al., 2009) and some gaming servers (Frey and Sumner, 2019). On reddit and most popular social media platforms, however access to moderation actions is difficult to obtain, and requires separate permission from the mods of every community to be studied (Li et al., 2022a, b). In contrast, by utilizing public discourse around moderators, our method works for every community whose content is public.
@@ 3. Methods
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al., 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, membership, and moderators. These attributes, along with the the fact that almost all content on reddit is publicly available, make it an ideal platform for studying community governance at a large scale.
In this work we make extensive use of publicly available reddit data from the Pushshift (Baumgartner et al., 2020). From these data, we detect and classify posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). In addition, we collect supplementary data about moderators (§3.1) and communities (§3.2).
@@ 3.1. Computing Moderator Timelines
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of the 10,000 largest subreddit’s moderators using reddit’s API and snapshots from the Wayback Machine, a web archiving service provided by the Internet Archive (int, [n. d.]). We start by scraping current moderator info pages for the 10,000 largest subreddits on the platform, directly from reddit. Each such page contains a list of the current moderators, in order of seniority, along with the exact timestamp that each moderator was added as a moderator to the subreddit. We then use the publicly accessible API from the Internet Archive’s Wayback Machine to scrape every archived copy of every subreddits’ moderator info page going back to 2010. We scrape 30,302 historical archived copies of moderator info pages, which, combined with the 10,000 present-day copies (one per subreddit), give us detailed timelines of who moderated each subreddit when, and when they started as a moderator. Due to the functionality of the Wayback Machine, we have higher temporal resolution (more archived snapshots) for larger and more popular subreddits, yet, because each moderator info page encodes the exact timestamp for each moderator’s start date, having fewer snapshots results only decreased accuracy of when moderators resign their posts, not when they are appointed — which is the primary focus of many of our analyses. Since the exact end time of a moderator’s tenure can only be inferred from examining when they were removed from the list of moderators for each subreddit, we adopt a conservative strategy which deliberately underestimates the length of moderators’ tenures: we consider the end date of their tenure as that of the last snapshot for which they were still listed as a moderator. Our method only misses moderators whose appointment, entire tenure, and resignation all occur within the ‘gap’ between snapshots [^2] We believe this is rare, as the average mod tenure we can detect is longer than four years, and the mean gap between snapshots is 56 days.. Note that to preserve moderator privacy, we do not make our moderator timelines public.
@@ 3.2. Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the topic of every subreddit included in our analyses, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld et al., 2022). The prompt used for this class is given in Appendix B.1. We used the manually-labeled dataset of 123 subreddits from Weld et al. (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld et al., 2022). We average across all survey responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
@@ 3.3. Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying posts and comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify content written by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects content discussing moderators (differentiating them from those which make other use of ‘mod’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of this content into positive, neutral, negative, and exclude classes. We only include content written by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not how mods discuss themselves in public. We apply our pipeline, which we make public, to all reddit posts and comments made from Jan. 2020 to June 2021.
@@
Prefilter Step Details (Step 1). The prefilter step efficiently identifies posts and comments which discuss moderators. We use a regular expression-based filter to find all posts and comments which include the words ‘mod(s),’ ‘moderator(s), and use our moderator tenure timelines to exclude posts and comments written by moderators during their appointment, as our goal is to measure how non-moderator community members discuss moderators, not how moderators discuss themselves.
Detection Step Details (Step 2). On reddit, the term ‘mods’ often refers to ‘moderators,’ but often is used as shorthand for ‘modifications,’ as in ‘video game mods’ or ‘car mods.’ To differentiate posts and comments which discuss moderators from those which use ‘mod(s)’ in other senses, we fine-tuned a RoBERTa-based binary classifier (Liu et al., 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the prefilter step results and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa (Figure 1). To label posts and comments for the classification task, two annotators worked together to iteratively refine a codebook (Appendix A), then independently labeled 200 randomly sampled posts/comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch, 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts/comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts/comments for use as a training set, of which 484 were randomly sampled, and 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa model trained for the classification task.
@@
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators (e.g., discussion of the /r/cats moderators taking place on /r/cats). For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with content that erroneously passed the detection step. We trained on this ‘exclude’ class in addition to those used for downstream analyses: positive, neutral, and negative.  [^3] Additionally, we attempted to classify specific complaints about moderation: excessive moderation, insufficient moderation, and biased moderation, but found classifier performance not to be amenable for high-confidence downstream analyses.
To further enhance the performance of our model, we performed data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al., 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al., 2023) fine-tuned with QLoRA (Dettmers et al., 2023), using the prompt shown in Appendix B.2. LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step left us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make an anonymized version of this dataset public.
@@
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4 & §6: The Amount of Mod Discourse for a given community is the fraction of all posts and comments discussing mods in that community, regardless of their sentiment. The Composition of Mod Discourse with positive/negative sentiment is the percent of all posts and comments discussing mods in a given community that have positive/negative sentiment. To ensure a meaningful amount of data, we exclude all communities that did not have at least one positive, neutral, and negative post or comment discussing moderators during our 18-month analysis period. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. This filtering leaves us with 5,282 subreddits used in downstream analyses.
3.4. Validating our Measures of Perceptions of Moderators
Our methods for detecting and classifying community members’ perceptions of moderators (§3.3) measure when community members publicly post or comment in their own communities to discuss moderators. For ethical as well as practical reasons, we cannot measure when community members discuss their moderators privately. However, there are two other instances where community members may attempt to discuss their moderators publicly but would not be included in our analyses. In this section, we show that these two cases are very rare relative to the public moderator discourse that we do measure.
For purposes of computational efficiency, our initial prefiltering step only includes posts and comments which use the phrase ‘mod(s)’ or ‘moderator(s)’ for subsequent classification. Occasionally, however, community members may discuss their moderators by mentioning their usernames specifically. For example, instead of writing ‘the mods of this subreddit are unfair,’ which would be captured by our pipeline, a community member might write ‘/u/exampleModerator is unfair.’ To assess the frequency of this form of moderator discourse, for a sample period of one month (January 2018, the first month of our analysis period), we compute how frequently the usernames of any active moderator is mentioned directly in their own subreddit, in addition to the frequency of uses of ‘mod(s)’ or ‘moderator(s).’ We find that such specific moderator mentions are very rare, with the average subreddit having 64.11 generic uses of ‘mod(s)’ or ‘moderator(s)’ for every time a moderator is mentioned specifically by their username. We assessed the impact that including these specific mentions would have on downstream analyses, and concluded qualitatively that they did not make a substantial difference in our results. Thus, for purposes of computational efficiency, we chose to exclude specific moderator mentions from our moderator discourse classification pipeline.
An additional potential source of bias in our analyses is that of removed posts and comments. On reddit, moderators are occasionally accused of removing content in their communities that is critical of the moderators (Matias, 2019a). As this content could be removed before entering our moderator discourse classification pipeline, it is conceivable that our results undercount moderator discourse in communities where the moderators remove such content. Assessing this source of bias requires knowledge of the content that is removed, which is technically infeasible to collect at a large scale. However, prior researchers have collected content before it is removed by moderators by scraping content from specific subreddits as it is posted, then checking later to see if the content has been removed (Chandrasekharan et al., 2018; Kumar et al., 2023). We used the Pushshift API (before it was taken offline) in the same manner, to collect the text of content that was removed from 100 randomly sampled subreddits over a two year period (2017-2018, inclusive). Using this method, we collected 263,657 pieces of removed content from 100 different communities. Of the content that was removed by moderators in this sample, we find that it is extremely rare for moderators to remove content that mentions the moderators: The average community in our sample removes 102 posts/comments which do not mention the moderators for every post or comment removed that does mention the moderators. Once again, we assessed the impact that including removed comments would have on our downstream analyses, and concluded that their inclusion would not make a substantial difference in our results (in addition to be infeasible at the scale of the rest of this work).
@@ 4. How are moderators of different communities perceived differently by their communities?
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities have more mod discourse (Figure 3a), with tiny communities with fewer than ten posts+comments per day having 7.6×7.6$\times$7.6 × as many posts and comments discussing moderators as huge communities (those with more than 10k posts and comments per day), relative to the total amount of content. Tiny communities also have 6.0×6.0$\times$6.0 × more positive mod discourse, and 0.46×0.46$\times$0.46 × as much negative discourse, than huge communities.
@@ 5. Adjusting for Confounders using IPTW & DID
While we are ideally interested in causal, actionable insights, our study is a retrospective observational study, and like any other, subject to potential confounding. In §4 we identified two important confounders, size and topic, which are correlated with perceptions of moderators (Figure 3). Next we describe two methods from the causal inference literature which we use to adjust for these and other confounding factors in subsequent analyses.
Inverse Probability of Treatment Weighting (IPTW). IPTW is a statistical method to adjust for imbalance in possible confounding factors when making comparisons between different observed groups (e.g., a treatment and a control group). We select IPTW over other methods, such as stratification, due to its excellent efficiency (Austin and Stuart, 2015). In §6, we compute the probability of treatment (propensity score, $P(Z|\mathbf{X})$) by applying a logistic regression model to community covariates $\mathbf{X}$ including the topic and size of the community. A complete list of covariates for each analysis is given in Appendix D. We generalize IPTW to non-binary treatments to estimate a dose-response curve using the method from (Althoff et al., 2016) in which the treatment is discretized into bins and a different propensity score is computed for each treatment bin in a ‘one-versus-rest’ scheme. For final analyses, each observation is weighted by the inverse of the probability of the treatment it received, such that the weight for observation $i$ is $w_{i}=\frac{Z_{i}}{P(Z_{i}=1|\mathbf{X}_{i})}+\frac{1-Z_{i}}{P(Z_{i}=0|\mathbf{X}_{i})}$, where $Z_{i}$ is the treatment received by observation $i$ (0 for control, 1 for treated), and $\mathbf{X}_{i}$ is a vector of the covariates. In analyses where we use IPTW, we also include the non-IPTW adjusted estimate in figures with a gray color. In these analyses, reweighting does not dramatically impact our findings, suggesting that the impact of these confounders is moderate at best.
An important validity check for IPTW is to assess the balance of covariates for each treatment group after weighting as applied (Austin and Stuart, 2015). Two groups are often considered ‘balanced’ or ‘indistinguishable‘ if all covariates are within a standardized mean difference (SMD) of 0.25 standard deviations (Althoff et al., 2016). For each treatment group, we compute the difference between each covariate’s weighted mean value and the reference distribution, consisting of the entire population. We compute the standardized mean difference (SMD) by normalizing the difference in means by the standard deviation of the values of the reference distribution. Our reweighting method achieves balance for every covariate substantially related to the treatment. SMDs after weighting are given for each covariate and each analysis in Appendix D.